i dont understand these guestions..can i please get some help
2. a) David works part-time for a bike repair shop. He earns $1 for each tire he installs and $4 for each gear mechanism he assembles. If he did a total of 50 insatallations and assemblies and earned $92, how many tires did he install?
b) A car rental angency rents cars under two plans:
Plan 1: $55 per day and $0.10/km
Plan 2: $40 per day and $0.25/km
Determine when both plans will result in the same charge for a one-day rental.

Answer by bam878s(77)

(a)
Let x = total number of tires he installs
and y = total number of gear mechanisms assembled.
We have,
x + y = 50 and
x + 4y = 92. If we subtract the first equation from the second equation we get
3y = 42. Now divide by 3 to solve for y
y = 14. So he assembled 14 gear mechanisms. To find out the number of tires we simply plug this value of y into one of the original two equations and solve for x
x + 14 = 50. Subtract 14 from both sides.
x = 36 tires he intalled.
(b)
Let x = number of km.
To solve for this one we simply set the equations equal to each other and solve for x.
55 + .10x = 40 + .25x
15 = .15x
divide by .15
100 = x.
So after 100 km, the charge will be the same.